A solenoid with a cross-sectional area of 1.75×10−3 m2 is 0.780 m long and has 450 turns per meter. Find the magnitude of induced emf in this solenoid if the current in it is increased from 0 to 1.65 A in 50.5 ms .
Answer #1

Induced emf is given by,

$$ \begin{aligned} e &=-L\left(\frac{d i}{d t}\right) \\ &=-\mu_{0} l A n^{2}\left(\frac{d i}{d t}\right) \\ &=-\left(4 \pi \times 10^{-7} \mathrm{H} / \mathrm{m}\right)(0.780 \mathrm{~m})\left(1.75 \times 10^{-3} \mathrm{~m}^{2}\right)(450 \text { turn } / \mathrm{m})^{2}\left(\frac{1.65 \mathrm{~A}}{50.5 \times 10^{-3} \mathrm{~s}}\right) \\ &=-0.0113 \mathrm{~V} \end{aligned} $$

Therefore, the induced emf is \(-0.0113 \mathrm{~V}\).